# A New Construction Structure on Multi-access Coded Caching with Linear Subpacketization: Cyclic Multi-Access Non-Half-Sum Disjoint Packing

### v1

@@ Abstract
We consider the $(K,L,M,N)$ multi-access coded caching system introduced by Hachem et al., which consists of a central server with $N$ files and $K$ cache nodes, each of memory size $M$, where each user can access $L$ cache nodes in a cyclic wrap-around fashion. At present, several existing schemes achieve competitive transmission performance, but their subpacketization levels grow exponentially with the number of users. In contrast, schemes with linear or polynomial subpacketization always incur higher transmission loads. We aim to design a multi-access coded caching scheme with linear subpacketization $F$ while maintaining low transmission load. Recently, Cheng et al. proposed a construction framework for coded caching schemes with linear subpacketization (i.e., $F=K$) called non-half-sum disjoint packing (NHSDP). Inspired by this structure, we introduce a novel combinatorial structure named cyclic multi-access non-half-sum disjoint packing (CMA-NHSDP) by extending NHSDP to MACC system. By constructing CMA-NHSDP, we obtain a new class of multi-access coded caching schemes. Theoretical and numerical analyses show that our scheme achieves lower transmission loads than some existing schemes with linear subpacketization. Moreover, the proposed schemes achieves lower transmission load compared to existing schemes with exponential subpacketization in some case.
@@ I-A Multi-access Coded Caching
Under the constraint of uncoded cache placement, the authors of [8] introduced a multi-access coded caching scheme. They proved that the gap between its achievable rate and the information-theoretic lower bound is no more than $cL$, where $c$ is a constant. Reference [22] propose a new scheme using index coding, which achieves a lower rate than [8] when $K<\frac{KLM}{N}+L$. Furthermore, for the case where $L\geq K/2$ and $N\geq K$, the authors of [22] derived a converse bound under uncoded placement. They proved that the scheme has order optimality within a factor of 2. A transformation method is proposed in [23] to adapt the MN scheme to the multi-access scenario. This adaptation achieves a lower rate than [8] when $L\nmid K$. The work in [24] decomposed the multi-access caching problem into structured index coding problem. Based on solutions to the structured index coding problem, a new scheme has been proposed with with a rate that does not exceed the rates of the schemes in [22] and [23]. It is worth noting that subpacketization of the above scheme increases exponentially with KK.[25] proposed a consecutive cyclic placement strategy. By constructing a class of PDAs under this strategy, a novel multi-access coded caching scheme was introduced. This scheme have a subpacketization increasing linearly with the number of users and attains a minimal rate of $\frac{K-KML/N}{2\left\lfloor\frac{K}{K-KML/N+1}\right\rfloor+1}$. All the aforementioned schemes are designed for general system parameters. A summary is provided in Table I. For specific parameter configurations, several schemes with reduced subpacketization have been introduced in [26, 27, 28, 29]. These are also included in Table I.
TABLE I: Known multi-access coded caching schemes with uncoded placement.Where $K$ is number of users, number of cache nodes is equal to $K$, $L$ is access degree and cache-node memory ratio is $\frac{t}{K}$. If the range of $t$ is not specified in the Limitation section, then $t\in\{0,\ldots,\lfloor\frac{K}{L}\rfloor\}$.
@@
In Table I, it can be observed that while the schemes in [8, 22, 23, 24] achieve competitive transmission rates, their subpacketization grow exponentially with the number of users. In contrast, [25, 28, 29] and other schemes maintain linear or polynomial subpacketization, but at the cost of limited coded caching gains and higher transmission loads. This trade-off reveals a key limitation in multi-access coded caching. Although the PDA serves as a combinatorial structure that characterizes placement and delivery strategies, its definition alone does not directly yield constructions that simultaneously achieve low subpacketization and significant coded caching gain. Therefore, constructing a general framework for PDA that guarantee linear subpacketization while maintaining near-optimal transmission performance is an important and meaningful research direction.
@@ I-B Research motivation and contribution
  • By introducing recursive offsets into the construction framework of NHSDP, we obtain a class of CMA-NHSDPs. Based on the CMA-NHSDPs, we can obtain a $(K=v,L,M,N)$ multi-access coded caching scheme with the memory ratio $M/N=\frac{v-2^{n}\prod_{i=1}^{n}m_{i}}{vL}$, coded caching gain $g=2^{n}$, and transmission load $R=\prod_{i=1}^{n}m_{i}$ for any odd integer $v\geq\prod_{i=1}^{n}(1+2m_{i})+2\sum_{i=1}^{n}\left(g(i)\prod_{j=i+1}^{n}(1+2m_{j})\right)$ where $n$ and $m_{1}$, $m_{2}$, $\ldots$, $m_{n}$ are any positive integers. In particular, when $\frac{m_{1}}{L}=m_{2}=\cdots=m_{n}=\left\lfloor\frac{(\frac{v}{L})^{1/n}-1}{2}\right\rfloor$, we obtain a $(K=v,L,M,N)$ multi-access coded caching scheme with the memory ratio $\frac{M}{N}=\frac{1}{L}-\frac{\left\lfloor(\frac{v}{L})^{\frac{1}{n}}-1\right\rfloor^{n}}{v}$, subpacketization $F=K$, coded caching gain $g=2^{n}$, and transmission load $R={\left\lfloor\frac{v^{1/n}-1}{2}\right\rfloor}^{n}L$.
@@ II-C Non-half-sum disjoint packing
By Lemma 2, in order to obtain a PDA with linear subpacketization the author in [19] introduce a construction of NHSDPs. The idea is to embed integers into a high-dimensional geometric space. Specifically, a subset $\mathcal{X}$ of $\mathbb{Z}_{v}$ is designed to ensure that each element in $\mathfrak{D}$ can be uniquely represented as a linear combination of elements in $\mathcal{X}$. Moreover, the half-sum of any two distinct elements within the same block can also be represented by $\mathcal{X}$, and the sets of coefficients for these two representations must be disjoint. $\square$
@@ III Cyclic Multi-access Non-half-sum disjoint packing
We can find that each column of $\mathbf{P}$ contains exactly $Z=v-bg=27-3\times 4=15$ stars, which occurs in each row as $L$ cyclic consecutive. So the conditions C22 and C44 hold. In addition, there are exactly $S=81$ vectors in $\mathbf{P}$, and each vector occurs at most once in each row and each column. i.e., the conditions C22 and C33-a of Definition 2 hold. Finally, let us consider condition C33-b. Let us first consider the entries $\mathbf{P}(5,0)=\mathbf{P}0,5)=(5,3)$. We have $\mathbf{P}(5,5)=\mathbf{P}(0,0)=*$ which satisfies condition C33-b. Similarly, we can check that all the vectors in $\mathbf{P}$ satisfy condition C33-b) of Definition 2. Therefore, $\mathbf{P}$ is an $L$-continuous $(27,27,15,81)$ PDA that can realize a MACC scheme with memory ratio $\frac{M}{N}=\frac{5}{27}$, subpacketization $F=K=27$, and load $R=3$.
@@ IV Constructions of PDAs via CMA-NHSDP
Under the selection in (16), the resulting PDA $\mathbf{Q}$ has the parameters
$K=q^{n}L,\ \ F=q^{n}L,\ \ Z=q^{n}L-\lfloor{q-1}\rfloor^{n}L,\ \ S=\lfloor\frac{q-1}{2}\rfloor^{n}q^{n}L^{2},$
which can realize a MACC scheme with the memory ratio $\frac{M}{N}=\frac{1}{L}-\frac{\lfloor{q-1}\rfloor^{n}}{q^{n}L}$ and the transmission load $R=\lfloor\frac{q-1}{2}\rfloor^{n}L$.
@@
When $q\geq 3$ in Theorem 4 is an odd integer, we have a $L$-continuous $(K=q^{n}L,F=q^{n}L,Z=q^{n}L-\lfloor{q-1}\rfloor^{n}L,S=\lfloor\frac{q-1}{2}\rfloor^{n}q^{n}L^{2})$ PDA which realizes a $F$-division $(K=v,L,M,N)$ MACC scheme with the memory ratio $\frac{M}{N}=\frac{1}{L}-\frac{\lfloor{q-1}\rfloor^{n}}{q^{n}L}$, subpacketization $F=q^{n}L$, and transmission load $R=\lfloor\frac{q-1}{2}\rfloor^{n}L$.
We now compare the performance of Theorem 4 and Construction 2 through a concrete example. Let $n=4$, $L=3$, and let $m_{2}$ vary in 1,2,3,4{1,2,3,4}. The corresponding parameters obtained from the two approaches are summarized in Fig. 2. It can be observed that Theorem 4 yields a smaller value of $v$ while maintaining the same objective function value $f$.
Figure 2: v tradeoff
@@ V Performance analysis
In this section, we will present theoretical and numerical comparisons with the existing schemes [22, 23, 25, 28, 8, 22, 23, 24, 29, 26, 28, 27, 30] respectively to show the performance of our new scheme in Theorem 4.
@@ V-A1 Comparison with the RK1 scheme in [22]
Let $K=q^{n}L$ and $i=q^{n}-\lfloor{q-1}\rfloor^{n}$ in [22], i.e., the $10$th row of Table I. We compare the RK1 scheme with the proposed scheme in the following cases:
$F_{\text{RK1}}$ $=\frac{K}{i}\cdot\binom{K-iL+i-1}{i-1}=\frac{q^{n}L}{q^{n}-\lfloor{q-1}\rfloor^{n}}\cdot\binom{q^{n}+\lfloor{q-1}\rfloor^{n}(L-1)-1}{q^{n}-\lfloor{q-1}\rfloor^{n}-1}$,
$R_{\text{RK1}}$ $=K(1-\frac{LM}{N})^{2}=\frac{\lfloor{q-1}\rfloor^{2n}L}{q^{n}}.$
@@
$\frac{F_{\text{RK1}}}{F}$ $=\frac{\frac{q^{n}L}{q^{n}-\lfloor{q-1}\rfloor^{n}}\cdot\binom{q^{n}+\lfloor{q-1}\rfloor^{n}(L-1)-1}{q^{n}-\lfloor{q-1}\rfloor^{n}-1}}{q^{n}L}=\frac{\binom{q^{n}+\lfloor{q-1}\rfloor^{n}(L-1)-1}{q^{n}-\lfloor{q-1}\rfloor^{n}-1}}{q^{n}-\lfloor{q-1}\rfloor^{n}}\geq\frac{q^{n}-\lfloor{q-1}\rfloor^{n}}{q^{n}-\lfloor{q-1}\rfloor^{n}}=1,$
$\frac{R_{\text{RK1}}}{R}$ $=\frac{\frac{\lfloor{q-1}\rfloor^{2n}L}{q^{n}}}{\lfloor\frac{q-1}{2}\rfloor^{n}L}=\frac{2^{n}\lfloor{q-1}\rfloor^{n}}{q^{n}}\approx(2-\frac{2}{q})^{n}$
As a result, compared to the RK1 scheme in [22], our scheme either obtaining smaller or same subpacketization. The multiplicative reduction amount by our transmission is $(2-\frac{2}{q})^{n}$. Only need $K\geq 3^{n}L$ ,the formula $(2-\frac{2}{q})^{n}\geq 1$ always holds.
@@ V-A2 Comparison with the CW scheme in [23]
Let $K=q^{n}L$ and $t=q^{n}-\lfloor{q-1}\rfloor^{n}$ in [23], i.e., the $10$th row of Table I. We compare the CW scheme with the proposed scheme in the following cases:
$F_{\text{CW}}$ $=K\cdot\binom{K-t(L-1)}{t}=q^{n}L\cdot\binom{q^{n}+\lfloor{q-1}\rfloor^{n}(L-1)}{q^{n}-\lfloor{q-1}\rfloor^{n}}$,
$R_{\text{CW}}$ $=\frac{K-tL}{t+1}=\frac{\lfloor{q-1}\rfloor^{n}L}{q^{n}-\lfloor{q-1}\rfloor^{n}+1}.$
@@
$\frac{F_{\text{CW}}}{F}$ $=\frac{q^{n}L\cdot\binom{q^{n}+\lfloor{q-1}\rfloor^{n}(L-1)}{q^{n}-\lfloor{q-1}\rfloor^{n}}}{q^{n}L}=\binom{q^{n}-\lfloor{q-1}\rfloor^{n}+\lfloor{q-1}\rfloor^{n}L}{q^{n}-\lfloor{q-1}\rfloor^{n}}$,
$\frac{R_{\text{CW}}}{R}$ $=\frac{\frac{\lfloor{q-1}\rfloor^{n}L}{q^{n}-\lfloor{q-1}\rfloor^{n}+1}}{\lfloor\frac{q-1}{2}\rfloor^{n}L}=\frac{2^{n}}{q^{n}-\lfloor{q-1}\rfloor^{n}+1}$
As a result, compared to the CW scheme in [23], our scheme either obtaining smaller subpacketization. The multiplicative reduction amount by our transmission is $\frac{2^{n}}{q^{n}-\lfloor{q-1}\rfloor^{n}+1}$. Only when $K=qL$ ,the formula $\frac{2^{n}}{q^{n}-\lfloor{q-1}\rfloor^{n}+1}=1$ always holds.

### v2

@@ Abstract
We consider the multi-access coded caching (MACC) system introduced by Hachem et al., which consists of a central server with $N$ files and $K$ cache nodes, each of memory size $M$, where each user can access $L$ cache nodes in a cyclic wrap-around fashion. At present, several existing schemes achieve competitive transmission performance, but their subpacketization levels grow exponentially with the number of users. In contrast, schemes with linear or polynomial subpacketization always incur higher transmission loads. We aim to design a multi-access coded caching scheme with linear subpacketization $F$ while maintaining low transmission load. Recently, Cheng et al. proposed a construction framework for coded caching schemes with linear subpacketization (i.e., $F=K$) called non-half-sum disjoint packing (NHSDP). Inspired by this structure, we introduce a novel combinatorial structure named cyclic multi-access non-half-sum disjoint packing (CMA-NHSDP) by extending NHSDP to MACC system. By constructing CMA-NHSDP, we obtain a new class of multi-access coded caching schemes. Theoretical and numerical analyses show that our scheme achieves lower transmission loads than some existing schemes with linear subpacketization. Moreover, the proposed scheme achievesa lower transmission load compared to existing schemes with exponential subpacketization in some cases.
@@ I-A Multi-access Coded Caching
Under the constraint of uncoded cache placement, the authors of [8] introduced a multi-access coded caching scheme. They proved that the gap between its achievable rate and the information-theoretic lower bound is no more than $cL$, where $c$ is a constant. Reference [22] proposes a new scheme using index coding, which achieves a lower rate than [8] when $K<\frac{KLM}{N}+L$. Furthermore, for the case where $L\geq K/2$ and $N\geq K$, the authors of [22] derived a converse bound under uncoded placement. They proved that the scheme has order optimality within a factor of 2. A transformation method is proposed in [23] to adapt the MN scheme to the multi-access scenario. This adaptation achieves a lower rate than [8] when $L\nmid K$. The work in [24] decomposed the multi-access caching problem into structured index coding problem. Based on solutions to the structured index coding problem, a new scheme has been proposed with with a rate that does not exceed the rates of the schemes in [22] and [23]. It is worth noting that subpacketization of the above scheme increases exponentially with KK.[25] proposed a consecutive cyclic placement strategy. By constructing a class of PDAs under this strategy, a novel multi-access coded caching scheme was introduced. This scheme have a subpacketization increasing linearly with the number of users and attains a minimal rate of $\frac{K-KML/N}{2\left\lfloor\frac{K}{K-KML/N+1}\right\rfloor+1}$. All the aforementioned schemes are designed for general system parameters. A summary is provided in Table I. For specific parameter configurations, several schemes with reduced subpacketization have been introduced in [26, 27, 28, 29]. These are also included in Table I.
TABLE I: Known multi-access coded caching schemes with uncoded placement.Where $K$ is number of users, number of cache nodes is equal to $K$, $L$ is the access degree and the cache-node memory ratio is $\frac{t}{K}$. If the range of $t$ is not specified in the Limitation section, then $t\in\{0,\ldots,\lfloor\frac{K}{L}\rfloor\}$.
@@
In Table I, it can be observed that while the schemes in [8, 22, 23, 24] achieve competitive transmission rates, their subpacketization grows exponentially with the number of users. In contrast, [25, 28, 29] and other schemes maintain linear or polynomial subpacketization, but at the cost of limited coded caching gains and higher transmission loads. This trade-off reveals a key limitation in multi-access coded caching. Although the PDA serves as a combinatorial structure that characterizes placement and delivery strategies, its definition alone does not directly yield constructions that simultaneously achieve low subpacketization and significant coded caching gain. Therefore, constructing a general framework for PDA that guarantee linear subpacketization while maintaining near-optimal transmission performance is an important and meaningful research direction.
@@ I-B Research motivation and contribution
  • We propose a CMA-NHSDP construction framework by introducing an offset function into the NHSDP construction. This offset function is designed to ensure that the differences between sorted integers within our proposed structure are $kL+1(k\in\mathbb{N})$, while maximizing the number of integers, i.e., minimizing the memory ratio. Based on the CMA-NHSDPs, we can obtain a $(K=v,L,M,N)$ multi-access coded caching scheme with the memory ratio $M/N=\frac{v-2^{n}\prod_{i=1}^{n}m_{i}}{vL}$, coded caching gain $g=2^{n}$, and transmission load $R=\prod_{i=1}^{n}m_{i}$ for any odd integer $v\geq\prod_{i=1}^{n}(1+2m_{i})+2\sum_{i=1}^{n}\left(g(i)\prod_{j=i+1}^{n}(1+2m_{j})\right)$ where $n$ and $m_{1}$, $m_{2}$, $\ldots$, $m_{n}$ are any positive integers. In particular, when $\frac{m_{1}}{L}=m_{2}=\cdots=m_{n}=\left\lfloor\frac{(\frac{v}{L})^{1/n}-1}{2}\right\rfloor$, we obtain a $(K=v,L,M,N)$ multi-access coded caching scheme with the memory ratio $\frac{M}{N}=\frac{1}{L}-\frac{\left\lfloor(\frac{v}{L})^{\frac{1}{n}}-1\right\rfloor^{n}}{v}$, subpacketization $F=K$, coded caching gain $g=2^{n}$, and transmission load $R={\left\lfloor\frac{v^{1/n}-1}{2}\right\rfloor}^{n}L$.
@@ II-C Non-half-sum disjoint packing
By Lemma 2, in order to obtain a PDA with linear subpacketization the authors in [19] introduce a construction of NHSDPs. The idea is to embed integers into a high-dimensional geometric space. Specifically, a subset $\mathcal{X}$ of $\mathbb{Z}_{v}$ is designed to ensure that each element in $\mathfrak{D}$ can be uniquely represented as a linear combination of elements in $\mathcal{X}$. Moreover, the half-sum of any two distinct elements within the same block can also be represented by $\mathcal{X}$, and the sets of coefficients for these two representations must be disjoint. $\square$
@@ III Cyclic Multi-access Non-half-sum disjoint packing
We can find that the stars occurs exactly $Z=v-bg=27-3\times 4=15$ times in each column of $\mathbf{P}$, and appears in each row as $L=3$ cyclic consecutive. So the conditions C22 and C44 hold. In addition, there are exactly $S=81$ vectors in $\mathbf{P}$, and each vector occurs at most once in each row and each column. i.e., the conditions C22 and C33-a of Definition 2 hold. Finally, let us consider condition C33-b. Let us first consider the entries $\mathbf{P}(5,0)=\mathbf{P}0,5)=(5,3)$. We have $\mathbf{P}(5,5)=\mathbf{P}(0,0)=*$ which satisfies condition C33-b. Similarly, we can check that all the vectors in $\mathbf{P}$ satisfy condition C33-b) of Definition 2. Therefore, $\mathbf{P}$ is an $L$-continuous $(27,27,15,81)$ PDA that can realize a MACC scheme with memory ratio $\frac{M}{N}=\frac{5}{27}$, subpacketization $F=K=27$, and load $R=3$.
@@ IV Constructions of PDAs via CMA-NHSDP
Under the selection in (16), the resulting $L$-continuous PDA has the parameters
$K=q^{n}L,\ \ F=q^{n}L,\ \ Z=q^{n}L-{\left(2\lfloor{\frac{q-1}{2}}\rfloor\right)}^{n}L,\ \ S=\lfloor\frac{q-1}{2}\rfloor^{n}q^{n}L^{2},$
which can realize a MACC scheme with the memory ratio $\frac{M}{N}=\frac{1}{L}-\frac{{\left(2\lfloor{\frac{q-1}{2}}\rfloor\right)}^{n}}{q^{n}L}$ and the transmission load $R=\lfloor\frac{q-1}{2}\rfloor^{n}L$.
@@
When $q\geq 3$ and $L$ in Theorem 4 is an odd integer, we have an $L$-continuous $(K=q^{n}L,F=q^{n}L,Z=q^{n}L-(q-1)^{n}L,S=(\frac{q-1}{2})^{n}q^{n}L^{2})$ PDA which realizes a $F$-division $(K=v,L,M,N)$ MACC scheme with the memory ratio $\frac{M}{N}=\frac{1}{L}-\frac{(q-1)^{n}}{q^{n}L}$, subpacketization $F=q^{n}L$, and transmission load $R=(\frac{q-1}{2})^{n}L$.
Next, we compare the performance of Theorem 4 and Construction 2 via a concrete example. Specifically, we set $n=4$, $L=3$, and let $m_{2}$ vary in 1,2,3,4{1,2,3,4}. The corresponding parameters derived from the two approaches are summarized in Fig. 2. It can be observed that Theorem 4 achieves a smaller value of $v$ while maintaining the same objective function value $f$.
Figure 2: $v$-$f$ tradeoff comparison between Theorem 4 and Construction 2
@@ V Performance analysis
In this section, we will present theoretical and numerical comparisons with the existing schemes [22, 25, 8, 23, 24, 28] respectively to show the performance of our new scheme in Theorem 4.
@@ V-A1 Comparison with the RK1 scheme in [22]
Let $K=q^{n}L$ and $t=q^{n}-2^{n}\lfloor\frac{q-1}{2}\rfloor^{n}$ in [22], i.e., the $10$th row of Table I. We compare the RK1 scheme with the proposed scheme in the following cases:
$F_{\text{RK1}}$ $=\frac{K}{t}\cdot\binom{K-tL+t-1}{t-1}=\frac{q^{n}L}{q^{n}-2^{n}\lfloor\frac{q-1}{2}\rfloor^{n}}\cdot\binom{q^{n}+2^{n}\lfloor\frac{q-1}{2}\rfloor^{n}(L-1)-1}{q^{n}-2^{n}\lfloor\frac{q-1}{2}\rfloor^{n}-1}$,
$R_{\text{RK1}}$ $=K(1-\frac{LM}{N})^{2}=\frac{2^{2n}\lfloor\frac{q-1}{2}\rfloor^{2n}L}{q^{n}}.$
@@
$\frac{F_{\text{RK1}}}{F}$ $=\frac{\frac{q^{n}L}{q^{n}-2^{n}\lfloor\frac{q-1}{2}\rfloor^{n}}\cdot\binom{q^{n}+2^{n}\lfloor\frac{q-1}{2}\rfloor^{n}(L-1)-1}{q^{n}-2^{n}\lfloor\frac{q-1}{2}\rfloor^{n}-1}}{q^{n}L}=\frac{\binom{q^{n}+2^{n}\lfloor\frac{q-1}{2}\rfloor^{n}(L-1)-1}{q^{n}-2^{n}\lfloor\frac{q-1}{2}\rfloor^{n}-1}}{q^{n}-2^{n}\lfloor\frac{q-1}{2}\rfloor^{n}}\geq\frac{q^{n}-2^{n}\lfloor\frac{q-1}{2}\rfloor^{n}}{q^{n}-2^{n}\lfloor\frac{q-1}{2}\rfloor^{n}}=1,$
$\frac{R_{\text{RK1}}}{R}$ $=\frac{\frac{2^{2n}\lfloor\frac{q-1}{2}\rfloor^{2n}L}{q^{n}}}{\lfloor\frac{q-1}{2}\rfloor^{n}L}=\frac{2^{2n}\lfloor\frac{q-1}{2}\rfloor^{n}}{q^{n}}\approx(2-\frac{2}{q})^{n}$
As a result, compared to the RK1 scheme in [22], our scheme either obtains smaller or same subpacketization. The multiplicative reduction amount by our transmission is $(2-\frac{2}{q})^{n}$. Only need $K\geq 3^{n}L$ ,the formula $(2-\frac{2}{q})^{n}\geq 1$ always holds.
@@ V-A2 Comparison with the CW scheme in [23]
Let $K=q^{n}L$ and $t=q^{n}-2^{n}\lfloor\frac{q-1}{2}\rfloor^{n}$ in [23], i.e., the $10$th row of Table I. We compare the CW scheme with the proposed scheme in the following cases:
$F_{\text{CW}}$ $=K\cdot\binom{K-t(L-1)}{t}=q^{n}L\cdot\binom{q^{n}+2^{n}\lfloor\frac{q-1}{2}\rfloor^{n}(L-1)}{q^{n}-2^{n}\lfloor\frac{q-1}{2}\rfloor^{n}}$,
$R_{\text{CW}}$ $=\frac{K-tL}{t+1}=\frac{2^{n}\lfloor\frac{q-1}{2}\rfloor^{n}L}{q^{n}-2^{n}\lfloor\frac{q-1}{2}\rfloor^{n}+1}.$
@@
$\frac{F_{\text{CW}}}{F}$ $=\frac{q^{n}L\cdot\binom{q^{n}+2^{n}\lfloor\frac{q-1}{2}\rfloor^{n}(L-1)}{q^{n}-2^{n}\lfloor\frac{q-1}{2}\rfloor^{n}}}{q^{n}L}=\binom{q^{n}-2^{n}\lfloor\frac{q-1}{2}\rfloor^{n}+2^{n}\lfloor\frac{q-1}{2}\rfloor^{n}L}{q^{n}-2^{n}\lfloor\frac{q-1}{2}\rfloor^{n}}$,
$\frac{R_{\text{CW}}}{R}$ $=\frac{\frac{2^{n}\lfloor\frac{q-1}{2}\rfloor^{n}L}{q^{n}-2^{n}\lfloor\frac{q-1}{2}\rfloor^{n}+1}}{\lfloor\frac{q-1}{2}\rfloor^{n}L}=\frac{2^{n}}{q^{n}-2^{n}\lfloor\frac{q-1}{2}\rfloor^{n}+1}$
As a result, compared to the CW scheme in [23], our scheme either obtains smaller subpacketization. The multiplicative reduction amount by our transmission is $\frac{2^{n}}{q^{n}-2^{n}\lfloor\frac{q-1}{2}\rfloor^{n}+1}$. Only when $K=qL$ ,the formula $\frac{2^{n}}{q^{n}-2^{n}\lfloor\frac{q-1}{2}\rfloor^{n}+1}=1$ always holds.